# PUCP-Metrix: A Comprehensive Open-Source Repository of Linguistic Metrics for Spanish

### v1

@@ Abstract
Linguistic features remain essential for interpretability and tasks involving style, structure, and readability, but existing Spanish tools offer limited coverage. We present PUCP-Metrix, an open-source repository of 182 linguistic metrics spanning lexical diversity, syntactic and semantic complexity, cohesion, psycholinguistics, and readability. PUCP-Metrix enables fine-grained, interpretable text analysis. We evaluate its usefulness on Automated Readability Assessment and Machine-Generated Text Detection, showing competitive performance compared to an existing repository and strong neural baselines. PUCP-Metrix offers a comprehensive, extensible resource for Spanish, supporting diverse NLP applications.
PUCP-Metrix: A Comprehensive Open-Source Repository of Linguistic Metrics for Spanish
@@ 1 Introduction
Linguistic features have gained renewed importance in explainable NLP, particularly for tasks requiring interpretability, stylistic sensitivity, or attention to surface-level properties. Despite advances in end-to-end neural models, recent work shows that handcrafted or derived features remain essential in applications such as AI-generated text detection (Kumarage et al., 2023; Ciccarelli et al., 2024; Petukhova et al., 2024), educational NLP (Mizumoto and Eguchi, 2023; Hou et al., 2025; Atkinson and Palma, 2025), and readability assessment (Zeng et al., 2024; Liu et al., 2025). In automated essay scoring, for instance, models incorporating linguistic features offer more transparent and pedagogically meaningful evaluations (Hou et al., 2025). These trends highlight the need for robust, modular repositories of linguistic metrics that can complement deep models.
Beyond NLP applications, these repositories also support linguistic research, offering standardized, quantifiable descriptions of texts across genres, registers, and proficiency levels (Jiang, 2016; Kuiken, 2023). They enable empirical analyses of morphosyntactic variation, cohesion, or lexical sophistication, and facilitate cross-linguistic comparisons.
Existing tools have demonstrated the value of this approach. For instance, Coh-Metrix (McNamara et al., 2010) provides extensive metrics for English across various linguistic levels. Similar resources include NILC-Metrix for Portuguese (Leal et al., 2023), Coh-Metrix-Esp for Spanish (Quispesaravia et al., 2016), and MultiAzterTest for Spanish, English, and Basque (Bengoetxea and Gonzalez-Dios, 2021).
In this work, we introduce PUCP-Metrix, a new open-source toolkit for extracting linguistic metrics from Spanish texts. It expands the range of available metrics across lexical, syntactic, discourse, psycholinguistic, and readability dimensions. In addition, we demonstrate its utility in two downstream tasks: Automated Readability Assessment and Machine-Generated Text Detection.
@@
  • PUCP-Metrix, a comprehensive and extensible open-source repository of linguistic metrics for Spanish, featuring metrics not available in existing resources. [^1] The code is available at https://github.com/iapucp/pucp-metrix.
@@
2 Related Work
Linguistic analysis tools have played a key role in understanding and quantifying text complexity. Coh-Metrix McNamara et al. (2010), a widely used tool for English, provides metrics capturing lexical, syntactic, semantic, and discourse characteristics of texts. These metrics support applications from educational assessment to psycholinguistic research, offering a detailed view of text complexity. Inspired by this framework, similar tools have been developed for other languages, adapting metrics to reflect language-specific features.
For Portuguese, NILC-Metrix Leal et al. (2023) provides a comprehensive set of over 200 metrics covering lexical, syntactic, semantic, discourse, and psycholinguistic dimensions, enabling detailed text analysis for educational and linguistic research. In Spanish, both Coh-Metrix-Esp Quispesaravia et al. (2016) and MultiAzterTest Bengoetxea and Gonzalez-Dios (2021) offer comparable capabilities. Coh-Metrix-Esp adapts the original Coh-Metrix to Spanish, implementing 45 linguistic features and is applied in a Automated Readability Assesment task. MultiAzterTest combines over 125 linguistic and stylistic features with machine learning classifiers to evaluate text complexity in Spanish, English, and Basque.
@@ 3 PUCP-Metrix
3.1 System Design
We used an open-source implementation of Coh-Metrix for the Spanish language Quispesaravia et al. (2016) as a starting point for our work. To implement new metrics, we analyzed the metrics in the tools described in Section 2 and consulted their implementation details when available.
@@ 3.2 Linguistic Categories and Metrics
For our processing needs, including tokenization, dependency parsing, and POS tagging, we utilized the NLP library spaCy. We developed custom pipelines to extract linguistic metrics from the texts efficiently. Following these steps, we compiled a collection of 182 linguistic metrics for Spanish texts. The complete list is available at Appendix A.
@@
  • Lexical Diversity: 22 indicators measure the diversity of the text’s vocabulary, including the type-token ratio for various word categories (nouns, verbs, etc.), noun density, verb density, adverb density, adjective density, the Maas Index Mass (1972), MLTD, and vocd McCarthy Philip M (2010). Our implementation extends these measures with type-token ratios for additional word categories and their lemmatized forms. Key indicators include the following:
@@
  • Psycholinguistics: 30 indicators, these reflect psycholinguistic properties of words, specifically how they are understood by humans: concreteness, imageability, familiarity, age of acquisition, valence and arousal. These psycholinguistic properties were collected from the EsPal database Duchon et al. (2013) and works from Stadthagen-Gonzalez et al. (2017):
@@
  • Word Information: 24 indicators, more detailed word-level statistics, such as: number of nouns, number of verbs, number of adverbs, number of adjectives and number of content words.
  • Referential Cohesion: 12 indicators, serve to measure the interconnections within a text: noun overlap, argument overlap, stem overlap, content word overlap and anaphor overlap.
  • Textual Simplicity: 4 indicators, measure the simplicity of the text using the ratio of short or large sentences, such as: proportion of short sentences, proportion of medium sentences, proportion of long sentences, proportion of very long sentences.
  • Semantic Cohesion: 8 indicators, assessing the degree of semantic relatedness between different parts of the text, such as: LSA overlap of adjacent sentences, LSA overlap of all sentences, LSA overlap of adjacent paragraphs.
  • Word Frequency: 16 indicators, various measurements involving the Zipf’s frequency for different kinds of words, such as rare nouns count, rare verbs count, rare adverbs count, rare content words count and mean word frequency.
@@ 3.3 Comparison with Existing Tools
Table 1 shows the number of linguistic metrics implemented in Coh-Metrix-Esp, MultiAzterTest and PUCP-Metrix (ours). PUCP-Metrix provides a broader coverage of linguistic metrics compared to CohMetrix-Esp and MultiAzterTest, comprising a total of 182 metrics across 13 categories. Notably, PUCP-Metrix includes metrics in categories that are entirely missing or underrepresented in the other tools, such as Semantic Cohesion, Textual Simplicity, and Psycholinguistics, with 8, 4, and 30 metrics, respectively. This way, PUCP-Metrix can capture higher-level discourse, cognitive readability, and psycholinguistic properties.
@@
Category CohMetrix-Esp MultiAzterTest PUCP-Metrix (ours)
@@ 4.1 Automated Readability Assessment (ARA)
We follow an approach similar to that of Vásquez-Rodríguez et al. (2022). The original work introduced a benchmark for ARA of Spanish texts. The authors combined multiple corpora labeled by language proficiency levels and proposed 2-label and 3-label classification schemes.
In contrast, our study comprises four publicly available datasets —CAES, Coh-Metrix-Esp, Kwiziq, and HablaCultura— to ensure reproducibility and open accessibility. We adopt the same label mappings described in the paper, adapting all texts to two readability classification schemas: 2-label (simple, complex) and 3-label (basic, intermediate, advanced). The dataset’s descriptions and the labeling strategy can be found in Appendix B.
@@
We implement a baseline model based on the Spanish variant of RoBERTa (named RoBERTa-BNE) Fandiño et al. (2022) [^3] The model was available at https://huggingface.co/PlanTL-GOB-ES/roberta-base-bne. We fine-tune RoBERTa-BNE for both the 2-label and 3-label classification tasks using the aforementioned splits.
@@ 4.2 Machine-Generated Text Detection
We adopt the AuTexTification 2023 shared task dataset Sarvazyan et al. (2023), which comprises over 160,000 texts in English and Spanish across five domains: tweets, reviews, news, legal, and how-to articles and generated by both human and large language models (machine).
@@
Following the shared task settings of AuTexTification and in line with the ARA task, we adopt RoBERTa-BNE as our baseline. It is fine-tuned on the official training splits and evaluated on the corresponding test splits to ensure comparability.
For both tasks, we trained various machine learning models: Logistic Regression (LR), XGBoost (XGB), Support Vector Machines (SVM) and Random Forest (RF) on the metrics extracted with both MultiAzter and PUCP-Metrix.
@@ 5 Results and Discussion
5.1 Automated Readability Assesment
Table 2 compares PUCP-Metrix, MultiAzter, and RoBERTa-BNE on two-label ARA/complexity classification. PUCP-Metrix slightly outperforms MultiAzter across simple and complex texts, achieving an overall F1 of 97.46 with XGBoost versus 97.24 for MultiAzter. In addition, XGBoost consistently yields the highest F1 scores, with Random Forest also performing competitively, while Logistic Regression and SVM score slightly lower.
RoBERTa-BNE achieves the best overall F1 of 98.30, indicating that although PUCP-Metrix captures rich linguistic cues, deep contextual models excel at detecting subtle semantic patterns.
@@ 5.1 Automated Readability Assesment
Table 3 compares PUCP-Metrix, MultiAzter, and RoBERTa-BNE on the 3-label ARA task. PUCP-Metrix again performs slightly better than Multiazter, reaching an overall F1 of 96.72 with XGBoost versus 96.56 for Multiazter, being XGBoost the model that achieves the highest scores.
Similarly to previous results, PUCP-Metrix does not surpass RoBERTa-BNE; that achieves the best overall F1 of 98.13, with near-perfect performance on Basic and Intermediate texts and strong results on Advanced ones.
@@ 5.2 Machine-Generated Text Detection
Table 4 shows the performance of various machine learning models using the metrics provided by PUCP-Metrix, and Multiazter. Also, it shows the performance of a RoBERTa-BNE model fine-tuned on the AuTexTification dataset and the overall performance of RoBERTa-BNE and the best model reported at the shared task.
In general, PUCP-Metrix consistently outperforms MultiAzter across classifiers. For human texts, PUCP-Metrix increases F1 scores from 42–51 (Multiazter) to 60–66, and for machine texts from 70–73 to 71–76, showing its ability to capture linguistic and structural cues critical for detecting human-written content. Tree-based models, particularly XGBoost and Random Forest, leverage PUCP-Metrix most effectively, achieving the highest overall F1 scores.
Compared to RoBERTa-BNE, PUCP-Metrix achieves more balanced performance across classes. While RoBERTa-BNE attains very high precision for human texts (93.96), its recall is low (37.86), yielding an F1 below PUCP-Metrix’s best result. This indicates that contextual embeddings may miss the diversity of human writing, whereas interpretable linguistic metrics maintain robust detection across both classes.
Furthermore, PUCP-Metrix slightly surpasses the best model reported in the shared task (F1 70.77), suggesting that integrating linguistic features with neural models could further improve classification performance.
Finally, we conduct an analysis about what are the linguistic metrics that are more important for classification. In general, Machine-generated text detection relies on features related to frequency, readability, and cohesion, while ARA tasks prioritize descriptive, syntactic, and simplicity features. Details are provided in Appendix C.
@@ B.2 Machine-generated Text Detection
Human-generated texts in AuTexTification were sourced from publicly available datasets, including MultiEURLEX Chalkidis et al. (2021) (legal), XLSUM/MLSUM Hasan et al. (2021); Scialom et al. (2020) (news), COAR/COAH Molina-González et al. (2014) (reviews), XLM-Tweets Barbieri et al. (2022) and TSD Leis et al. (2019) (tweets), and WikiLingua Ladhak et al. (2020) (how-to articles). Machine-generated texts were produced using six large language models: three from the BLOOM family (BLOOM-1B7777Available at https://huggingface.co/bigscience/bloom-1b7., BLOOM-3B [^8] Available at https://huggingface.co/bigscience/bloom-3b., BLOOM-7B1999Available at https://huggingface.co/bigscience/bloom-7b1.) and three from the GPT-3 family (babbage, curie, text-davinci-003).

### v2

# PUCP-Metrix: An Open-source and Comprehensive Toolkit for Linguistic Analysis of Spanish Texts

@@ Abstract
Linguistic features remain essential for interpretability and tasks that involve style, structure, and readability, but existing Spanish tools offer limited coverage. We present PUCP-Metrix, an open-source and comprehensive toolkit for linguistic analysis of Spanish texts. PUCP-Metrix includes 182 linguistic metrics spanning lexical diversity, syntactic and semantic complexity, cohesion, psycholinguistics, and readability. It enables fine-grained, interpretable text analysis. We evaluate its usefulness on Automated Readability Assessment and Machine-Generated Text Detection, showing competitive performance compared to an existing repository and strong neural baselines. PUCP-Metrix offers a comprehensive and extensible resource for Spanish, supporting diverse NLP applications.
PUCP-Metrix: An Open-source and Comprehensive Toolkit for Linguistic Analysis of Spanish Texts
@@ 1 Introduction
Linguistic features have gained renewed importance in explainable NLP, particularly for tasks requiring interpretability, stylistic sensitivity, or attention to surface-level properties. Despite advances in end-to-end neural models, recent work shows that handcrafted or derived features remain essential in applications such as AI-generated text detection (Kumarage et al., 2023; Ciccarelli et al., 2024; Petukhova et al., 2024), educational NLP (Mizumoto and Eguchi, 2023; Hou et al., 2025; Atkinson and Palma, 2025), and readability assessment (Zeng et al., 2024; Liu et al., 2025). In automated essay scoring, for instance, models incorporating linguistic features offer more transparent and pedagogically meaningful evaluations (Hou et al., 2025). These trends highlight the need for robust, modular repositories/toolkits that allow to extract linguistic metrics that complement deep models.
Beyond NLP applications, these toolkits also support linguistic research, offering standardized, quantifiable descriptions of texts across genres, registers, and proficiency levels (Jiang, 2016; Kuiken, 2023). They enable empirical analyses of morphosyntactic variation, cohesion GonzalezDiosBengoetxea2021_VaxxStance, or lexical sophistication CROSSLEY201846, and facilitate cross-linguistic studies Ucar2024.
Existing tools have demonstrated the value of this approach. For instance, Coh-Metrix (McNamara et al., 2010) provides extensive metrics for English across various linguistic levels. Similar resources include NILC-Metrix for Portuguese (Leal et al., 2023), Coh-Metrix-Esp for Spanish (Quispesaravia et al., 2016), and MultiAzterTest for Spanish, English, and Basque (Bengoetxea and Gonzalez-Dios, 2021). Nevertheless, the latter two exhibit limited metric coverage and present challenges related to installing and inference efficiency.
In this work, we introduce PUCP-Metrix, a new open-source toolkit for extracting linguistic metrics from Spanish texts. It extends coverage to lexical, syntactic, discourse, psycholinguistic, and readability dimensions, providing a total of 182 linguistic metrics. In addition, we demonstrate its utility in two downstream tasks: Automated Readability Assessment and Machine-Generated Text Detection.
@@
  • PUCP-Metrix, a comprehensive and extensible open-source toolkit of linguistic metrics for Spanish, featuring metrics not available in existing resources. [^1] The code is available at https://github.com/iapucp/pucp-metrix.
@@ 2 PUCP-Metrix
PUCP-Metrix is a modular and extensible linguistic analysis toolkit for Spanish, designed to support both research and large-scale text processing through a Python library. Its architecture emphasizes flexibility and scalability, enabling users to efficiently extract a wide range of linguistic metrics from texts.
To achieve this, we leveraged the widely adopted NLP library spaCy for core processing tasks such as tokenization, part-of-speech tagging, and dependency parsing. Building on spaCy’s modular architecture, we developed custom pipelines that include both general-purpose and category-specific metrics implemented as reusable components. This design allows external users to easily extend or modify the system, ensuring that PUCP-Metrix remains both efficient and adaptable to new linguistic analyses.
@@ 2.1 Linguistic Categories and Metrics
We employed an open-source Spanish implementation of Coh-Metrix Quispesaravia et al. (2016) to collect initial linguistic metrics and guide our design. To develop additional metrics, we examined the implementations provided by tools such as MultiAzterTest and NILC-Metrix. Overall, we compiled 182 linguistic metrics for Spanish texts. The complete list is available at Appendix A.
@@
  • Lexical Diversity: 22 indicators measure the diversity of the text’s vocabulary, including the type-token ratio for various word categories (nouns, verbs, etc.), noun density, verb density, adverb density, and adjective density. Our implementation extends these measures with type-token ratios for additional word categories and their lemmatized forms. Another key indicators include the following:
@@
    • Brunet Index: A readability measure of lexical richness, where lower values indicate greater vocabulary diversity.
@@
  • Psycholinguistics: 30 indicators, reflecting psycholinguistic properties of words, specifically how they are understood by humans: concreteness, imageability, familiarity, age of acquisition, valence and arousal. These psycholinguistic properties were collected from the EsPal database Duchon et al. (2013) and works from Stadthagen-Gonzalez et al. (2017):
@@
  • Word Information: 24 indicators with more detailed word-level statistics, such as: number of nouns, number of verbs, number of adverbs, number of adjectives and number of content words.
  • Referential Cohesion: 12 indicators that serve to measure the interconnections within a text: noun overlap, argument overlap, stem overlap, content word overlap and anaphor overlap.
  • Textual Simplicity: 4 indicators, measuring the simplicity of the text using the ratio of short or large sentences, such as: proportion of short sentences, proportion of medium sentences, proportion of long sentences, proportion of very long sentences.
  • Semantic Cohesion: 8 indicators, assessing the degree of semantic relatedness between different parts of the text, such as: Latent Semantic Analysis (LSA) overlap of adjacent sentences, LSA overlap of all sentences, LSA overlap of adjacent paragraphs.
  • Word Frequency: 16 indicators, various measurements involving the Zipf’s frequency [^2] Zipf’s frequency is estimate by using the wordfreq tool. It is available at https://github.com/rspeer/wordfreq/. for different kinds of words, such as rare nouns count, rare verbs count, rare adverbs count, rare content words count [^3] Rare words were defined in a similar way as Bengoetxea and Gonzalez-Dios (2021). and mean word frequency.
@@ 2.2 Comparison with Existing Tools
Table 1 compares Coh-Metrix-Esp, MultiAzterTest, and PUCP-Metrix across 3 dimensions: ease of installation, inference time (i.e., the time required to extract linguistic features), and metric coverage. PUCP-Metrix improves upon existing tools by combining straightforward installation, fast inference, and a broad set of linguistic metrics. In contrast, MultiAzterTest, while providing many metrics, is challenging to install and depends on external tools such as syntactic parsers, resulting in slower feature extraction and limited scalability. PUCP-Metrix, by offering 182 metrics—significantly more than Coh-Metrix-Esp (48) and MultiAzterTest (141)—provides a practical and comprehensive solution for research in computational linguistics.
Easy to install Inference Time Metric Coverage
Coh-Metrix-Esp Fast 48
MultiAzterTest Slow 141
PUCP-Metrix (ours) Fast 182
Table 1: Comparison of PUCP-Metrix with existing tools for Spanish.
Table 2 shows the number of linguistic metrics implemented in Coh-Metrix-Esp, MultiAzterTest and PUCP-Metrix (ours). PUCP-Metrix provides a broader coverage of linguistic metrics compared to Coh-Metrix-Esp and MultiAzterTest, across 13 categories. Notably, PUCP-Metrix includes metrics in categories that are entirely missing or underrepresented in the other tools, such as semantic cohesion, textual simplicity, and psycholinguistics. This way, PUCP-Metrix can capture higher-level discourse, cognitive readability, and psycholinguistic properties.
@@
Category Coh-Metrix-Esp MultiAzterTest PUCP-Metrix (ours)
@@ 3.1 Automated Readability Assessment (ARA)
We adopt an approach similar to that of Vásquez-Rodríguez et al. (2022), who introduced a benchmark for ARA on Spanish texts. Their work unified both public and non-public corpora annotated with language proficiency levels and proposed two- and three-class classification schemas.
In contrast, our study comprises only four publicly available datasets —CAES, Coh-Metrix-Esp, Kwiziq, and HablaCultura— to ensure reproducibility and open accessibility. We adopt the same label mappings described in the paper, adapting all texts to two readability classification schemas: 2-label (simple, complex) and 3-label (basic, intermediate, advanced). The dataset’s descriptions and the labeling strategy can be found in Appendix B.
@@ 3.2 Machine-Generated Text Detection
We adopt the AuTexTification 2023 shared task dataset Sarvazyan et al. (2023), which comprises over 160,000 texts in English and Spanish across five domains: tweets, reviews, news, legal, and how-to articles generated by both human and large language models.
@@
3.3 Models
For both tasks, we trained several machine learning models—Logistic Regression (LR), XGBoost (XGB), Support Vector Machines (SVM), and Random Forests (RF)—using the metrics extracted with both MultiAzterTest and PUCP-Metrix.
Following the AuTexTification shared task setup and consistent with the ARA formulation, we use a RoBERTa-based model Fandiño et al. (2022) (RoBERTa-BNE) [^5] Model available at https://huggingface.co/PlanTL-GOB-ES/roberta-base-bne as our baseline. We fine-tune this model on the official training splits and evaluate it on the corresponding test splits to ensure comparability. Similarly, we fine-tune RoBERTa-BNE on both the 2-label and 3-label ARA classification schemas.
@@ 4 Results and Discussion
4.1 Automated Readability Assessment
Table 3 reports the results for the 2-label ARA task. PUCP-Metrix slightly outperforms MultiAzter, achieving an overall F1 of 97.46 with XGBoost compared to 97.24. However, this difference is not significant. XGBoost consistently yields the highest scores, followed by Random Forest, while Logistic Regression and SVM lag behind. RoBERTa-BNE achieves the best overall F1 (98.30), indicating that deep contextual models capture subtle semantic patterns beyond what feature-based metrics provide.
@@ 4.1 Automated Readability Assessment
Table 4 shows the 3-label ARA results. PUCP-Metrix again slightly surpasses MultiAzter (F1 96.72 vs. 96.56), with XGBoost as the top-performing classifier. RoBERTa-BNE remains the strongest model, achieving an overall F1 of 98.13 and near-perfect performance on Basic and Intermediate texts.
@@ 4.2 Machine-Generated Text Detection
Table 5 reports the performance of machine learning models using PUCP-Metrix and MultiAzter metrics, alongside a RoBERTa-BNE model fine-tuned on AuTexTification and the shared task’s best-reported results.
PUCP-Metrix consistently outperforms MultiAzter. For human texts, F1 increases from 42–51 (MultiAzter) to 60–66, and for machine texts from 70–73 to 71–76, showing its ability to capture linguistic and structural cues. Tree-based models, especially XGBoost and Random Forest, benefit most, achieving the highest overall F1.
Compared to RoBERTa-BNE, PUCP-Metrix provides more balanced class performance. While RoBERTa-BNE attains high precision on human texts (86.28), low recall (46.87) limits its F1, suggesting that contextual embeddings may miss the diversity of human writing. PUCP-Metrix also slightly surpasses the shared task’s best-reported F1 (70.77), indicating that integrating linguistic features with neural models could further improve results.
Finally, we analyze which linguistic metrics contribute most to classification. Overall, detecting machine-generated text depends primarily on features related to frequency, readability, and cohesion, whereas ARA tasks are driven by descriptive, syntactic, and simplicity-related features. Full details of the feature analysis are provided in Appendix C.
@@ B.2 Machine-generated Text Detection
Human-generated texts in AuTexTification were sourced from publicly available datasets, including MultiEURLEX Chalkidis et al. (2021) (legal), XLSUM/MLSUM Hasan et al. (2021); Scialom et al. (2020) (news), COAR/COAH Molina-González et al. (2014) (reviews), XLM-Tweets Barbieri et al. (2022) and TSD Leis et al. (2019) (tweets), and WikiLingua Ladhak et al. (2020) (how-to articles). Machine-generated texts were produced using six large language models: three from the BLOOM family (BLOOM-1B7999Available at https://huggingface.co/bigscience/bloom-1b7., BLOOM-3B [^10] Available at https://huggingface.co/bigscience/bloom-3b., BLOOM-7B1111111Available at https://huggingface.co/bigscience/bloom-7b1.) and three from the GPT-3 family (babbage, curie, text-davinci-003).